You are on the surface of an unknown planet and you wish to know what objects weigh here. When you push a small box sitting on a frictionless surface in the horizontal direction with a force of F 10.0 N, it moves 20.0 m in the first 2.0 s. Next, you release the box from rest from a distance of 10.0 m off the ground and it takes 2.50 seconds to reach the ground. What does the box weigh on the unknown planet?
Answer #1

here,

for the horizontal direction

let the mass be m

force appllied , F = 10 N

the accelration , a = F/m = 10/m

time taken , t = 2 s

s = 20 m

using seccond equation of motion

s = 0 + 0.5 * a * t^2

20 = 0 + 0.5 * 10/m * 2^2

m = 1 kg

for vertical direction

h = 10 m

let the gravitational accelration be g

using seccond equation of motion

h = 0 + 0.5 * g * t^2

10 = 0 + 0.5 * g * 2.5^2

solving for g

g = 3.2 m/s^2

the weight of box , W = m * g

W = 1 * 3.2 N = 3.2 N